# TiC-LM: A Web-Scale Benchmark for Time-Continual LLM Pretraining

### v2

@@ 1 Introduction
Figure 1: Replay allows for matching repeated from-scratch training. Each line traces a training trajectory for a continual learning run where the annotations indicate the data cutoff year of the checkpoint (e.g., “24” = 2024). We find that combining autoregressive (AR) learning rate schedules and data replay (red) can nearly match the perplexity on all months achieved by the Oracle series which re-trains from scratch every two years (gray), despite requiring 2.6$\times$ less compute. Meanwhile, methods that only modify the optimizer (blue) or loss (green) insufficiently prevent forgetting and plateau.
@@ 3 TiC-CC: over 10 years of web data
Data processing. We build upon the assets from DataComp-LM (Li et al., 2024a), starting with DCLM-Pool (Li et al., 2024a), which contains all CC dumps up to Dec-2022 and parsed to extract plaintext from webpages via the resiliparse library (Bevendorff et al., 2018, 2021). We split this data by month and reuse the same download and processing scripts to extend DCLM-Pool until July-2024. Next, we follow DCLM-Baseline’s pipeline by applying heuristic filters from RefinedWeb (Penedo et al., 2023) and a fuzzy-deduplication step which we modify to run only within each month rather than across all months. Alternatively, as in TiC-CLIP, one could deduplicate newer months against older ones. Here, we do not do this by default given that it may not always be helpful (e.g., it may remove Wikipedia pages where a few key facts have changed but most text remains the same). Instead, we allow for exploring the benefits/pitfalls of such data-centric interventions as part of method design. Finally, we do not use the final classifier-based filter in DCLM-Baseline, as this classifier was trained on data from all months, violating causality. For more details about the data pipeline, see Appx. A.
@@ 5 Continual Learning Baseline Methods
  • Cyclic Cosine is the simplest alternative which applies cosine decay within each training month using the same maximum learning rate and warmup. This was found to be most effective in TiC-CLIP (Garg et al., 2024).
  • Cyclic Cosine + AR (autoregressive) uses Cyclic Cosine except the maximum learning rate in each cycle decays across rounds, regressed from a global cosine schedule. It was shown to offer improvements by Roth et al. (2024).
@@
Non-continual Oracles. The alternative to training in a continual fashion is to simply retrain a model from scratch on an equal number of tokens from all currently available months. We refer to such a model as Oracle-$t$ where $t$ corresponds to a particular cutoff date. We then consider as a baseline a series of Oracle models re-trained roughly every two years (i.e., {2013-05, 2015-01, 2017-01, 2019-01, 2021-01, 2023-01, 2024-07}), where for any timestamped evaluation, the most recently trained Oracle is always used.
@@ 6 Experiments
Figure 3: Continual methods incur different trade-offs between ID and Backward performance on TiC-CC. We plot the regret matrices for selected methods at 440B token scale (lower is better). Overall, Cyclic Cosine leads to strong ID performance (along the diagonal) but also significant forgetting after a few years. This can be partially addressed by using an AR meta-schedule and more significantly by replay. However, too much replay such as with an $\alpha_{t}=1/t$ schedule scales poorly with a large number of timestamps, significantly sacrificing ID performance.
Table 2: Loss-based evaluations for various methods. We report log-perplexity values relative to Oracle-2024-07. While various optimizer (top) and regularization-based (bottom) methods trade-off backward transfer with ID performance, replay (middle) is required to obtain the least amount of forgetting. Bold values are within one standard deviation of the best in each column for a given token budget, with standard deviations estimated from three runs of Cyclic Cosine. Highlighted values indicate when a continual method outperforms the Oracle series.
Method Tokens TiC-CC $\downarrow$ TiC-CC-Wiki $\downarrow$ TiC-CC-News $\downarrow$
@@
Cyclic Cosine (std) 220B 0.072 0.027 0.161 0.038 0.032 0.074 0.058 0.015 0.109
Cyclic Cosine + AR 220B 0.058 0.040 0.166 0.032 0.031 0.074 0.041 0.017 0.110
Cyclic Rsqrt 220B 0.065 0.030 0.162 0.033 0.030 0.073 0.049 0.015 0.108
ScheduleFree 220B 0.065 0.036 0.164 0.036 0.033 0.076 0.049 0.017 0.110
Replay ($\alpha=1/t$) 220B 0.023 0.074 0.178 0.020 0.036 0.078 0.005 0.035 0.117
Replay ($\alpha_{t}=1/2$) 220B 0.024 0.042 0.167 0.024 0.031 0.074 0.013 0.019 0.111
Replay ($\alpha=1/t$) + AR 220B 0.026 0.083 0.181 0.019 0.037 0.079 0.004 0.039 0.119
Replay ($\alpha=1/2$) + AR 220B 0.025 0.055 0.171 0.022 0.032 0.076 0.009 0.022 0.112
LwF 220B 0.072 0.027 0.161 0.038 0.032 0.074 0.058 0.015 0.109
EWC 220B 0.061 0.032 0.162 0.031 0.029 0.071 0.046 0.014 0.108
Cyclic Cosine (std) 440B 0.082 -0.011 0.145 0.029 0.015 0.059 0.071 -0.001 0.099
Cyclic Cosine + AR 440B 0.058 -0.002 0.148 0.014 0.009 0.057 0.044 -0.005 0.097
Cyclic Rsqrt 440B 0.067 -0.007 0.146 0.018 0.010 0.057 0.055 -0.004 0.097
Schedule-Free 440B 0.063 -0.004 0.147 0.017 0.011 0.059 0.049 -0.004 0.098
Replay ($\alpha=1/t$) 440B 0.001 0.044 0.164 0.003 0.016 0.062 -0.009 0.013 0.105
Replay ($\alpha=1/2$) 440B 0.007 0.007 0.151 0.010 0.013 0.059 0.005 -0.000 0.099
Replay ($\alpha=1/t$) + AR 440B -0.003 0.050 0.166 -0.006 0.013 0.061 -0.017 0.014 0.105
Replay ($\alpha=1/2$) + AR 440B -0.002 0.016 0.154 -0.001 0.009 0.057 -0.009 -0.002 0.098
LwF 440B 0.082 -0.011 0.145 0.029 0.015 0.059 0.072 -0.001 0.099
EWC 440B 0.055 0.011 0.152 0.017 0.015 0.061 0.041 0.001 0.100
Oracle Series 1.16T -0.003 0.037 0.163 0.004 0.017 0.066 -0.007 0.020 0.107
@@
Training details. We train 3B parameter models using OpenLM (Gururangan et al., 2023). Each method observes a fixed total number of 220B or 440B tokens, equivalent to 4$\times$ and 8$\times$ the Chinchilla optimal amount.  [^2] Here, token counts are given by 20 $\times$ parameters $\times$ Chinchilla multiplier with a 1$\times$ multiplier being a near-optimal compute allocation found by Hoffmann et al. (2022). For continual runs, we further assume that current practitioners are (a) likely to have access to more than enough data to train initial models; (b) unlikely to wait to obtain non-trivial performance. Hence, we front-load the total token budget such that 110B is allocated to an initial pretraining on the first month (May-2013). Then, the remaining tokens are split equally among the other 113 continual timesteps. For our Oracle-$t$ runs, each is trained on the same number of tokens that a 220B continual run observes by the end of month $t$, totaling 1.16T tokens for all seven Oracle models together. Finally, to perform realistic hyperparameter selection, we follow takeaways from (Cha and Cho, 2024) and only use the first 10 timesteps for tuning (see Appx. C for more details).
Evaluation metrics. Each run produces a $T_{t}\times T_{e}$ matrix of evaluations $E$ where $T_{t},T_{e}$ are the total number of training/evaluation timesteps, $E_{i,j}$ is the performance of the model trained after seeing data up to month $i$ and evaluated on the month $j$.
@@
To control for inherent difficulty differences across evaluation months, we measure the regret $R_{i,j}=E_{i,j}-E^{*}_{j}$ where $E^{*}_{j}$ is the performance of Oracle-2024-07 on month $j$. Following Garg et al. (2024), we consider the following summary metrics, first defined assuming $T_{t}=T_{e}=T$ (deferring the discussion of the $T_{t}\neq T_{e}$ case to Sec. B.4):
@@
6.1 Held-out performance on TiC-CC
In Tab. 2, we first explore how well continual methods learn and retain various parts of our general web-scale training distribution, TiC-CC. Overall, we observe significant trade-offs between ID and backward transfer, with our findings as follows:
@@ 6.1 Held-out performance on TiC-CC
Replay is essential for addressing forgetting. Based on backward transfer and Figs. 16 and 15, all non-replay methods show significant forgetting at later checkpoints. Optimizer tweaks and EWC can somewhat reduce forgetting by sacrificing some ID performance. However, additionally applying replay can further improve backward transfer by 60% for 220B runs, a gap which only widens as we scale up to 440B tokens. Between replay variants, $\alpha_{t}=1/t$ achieves the lowest forgetting but $\alpha_{t}=1/2$ offers a better practical trade-off on TiC-CC, resulting in marginally worse Backward but substantially better ID. This differs from TiC-CLIP’s recommendation of $\alpha_{t}=1/t$, likely since decreasing the ratio of new data becomes problematic in our setup which uses 10$\times$ more timesteps.
@@ 6.2 Downstream evaluations
Continual methods outperform Oracle re-training on many downstream evaluations. While Replay ($\alpha_{t}=1/2$) + AR at 440B tokens outperforms the Oracle series on TiC-CC, the comparison is more nuanced for downstream tasks. All continual methods show significant gains on TiC-Wiki evaluations, while the Oracle series maintains an edge on TiC-CodeDocs-NumPy and TiC-StackE-English (see Appx. D). Meanwhile, methods like Cyclic Cosine outperform the Oracles on TiC-StackOverflow and TiC-CodeDocs-PyTorch but fall short on TiC-StackE-Math where only Replay $(\alpha=1/t)$ can do so.
Table 3: Benchmarking continual methods on downstream evaluations. For all dynamic evaluations, we report perplexity values relative to Oracle-2024-07 with log-scaling. Meanwhile, Core is an average of the accuracies of 22 downstream zero/few-shot tasks used by Li et al. (2024a), evaluated only on the final model checkpoint (without scaling by an Oracle). Bold values are within one standard deviation (estimated with 3 runs of Cyclic Cosine) of the best in each column. Highlighted values indicate matching the Oracle series (within one standard deviation).
[…]
Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd.
Cyclic Cosine (std) 440B 0.018 0.035 0.073 0.025 0.034 0.058 0.017 0.053 0.142 -0.017 -0.045 0.175
Cyclic Cosine + AR 440B 0.009 0.033 0.071 0.010 0.028 0.054 0.026 0.064 0.142 0.015 -0.017 0.195
Cyclic Rsqrt 440B 0.012 0.032 0.070 0.014 0.030 0.053 0.024 0.062 0.145 0.003 -0.027 0.187
ScheduleFree 440B 0.012 0.035 0.073 0.014 0.032 0.056 0.038 0.072 0.144 0.032 0.010 0.197
Replay ($\alpha=1/t$) 440B 0.019 0.045 0.078 0.014 0.034 0.057 0.039 0.090 0.176 0.109 0.086 0.238
Replay ($\alpha=1/2$) 440B 0.015 0.037 0.073 0.017 0.032 0.055 0.027 0.072 0.158 0.022 -0.002 0.189
Replay ($\alpha=1/t$) + AR 440B 0.015 0.044 0.077 0.005 0.031 0.056 0.057 0.103 0.179 0.143 0.122 0.261
Replay ($\alpha=1/2$) + AR 440B 0.010 0.036 0.073 0.007 0.029 0.054 0.036 0.079 0.160 0.060 0.031 0.214
LwF 440B 0.019 0.034 0.072 0.026 0.034 0.056 0.014 0.054 0.142 -0.025 -0.048 0.171
EWC 440B 0.015 0.039 0.074 0.016 0.036 0.058 0.039 0.077 0.155 0.058 0.028 0.211
Oracle Series 1.16T 0.021 0.052 0.082 0.018 0.042 0.062 0.023 0.074 0.151 0.047 0.000 0.218
Bwd. ID Fwd. Bwd. ID Fwd. Core (DCLM)
Cyclic Cosine 440B 0.049 0.030 0.012 0.053 0.109 0.065 49.6
Cyclic Cosine + AR 440B 0.023 0.009 -0.001 0.044 0.084 0.055 48.8
Cyclic Rsqrt 440B 0.033 0.017 0.005 0.049 0.095 0.060 49.4
ScheduleFree 440B 0.036 0.022 0.009 0.081 0.105 0.072 49.4
Replay ($\alpha=1/t$) 440B -0.025 -0.026 -0.022 0.036 0.026 0.041 49.4
Replay ($\alpha=1/2$) 440B $0$0.008 -0.011 0.029 0.049 0.044 50.1
Replay ($\alpha=1/t$) + AR 440B -0.028 -0.027 -0.022 0.032 0.025 0.047 49.3
Replay ($\alpha=1/2$) + AR 440B -0.015 -0.019 -0.017 0.036 0.042 0.043 49.3
LwF 440B 0.049 0.023 0.013 0.058 0.121 0.070 49.6
EWC 440B 0.014 0.004 -0.001 0.063 0.075 0.059 49.0
Oracle Series 1.16T -0.024 -0.028 -0.021 0.012 0.013 0.017 50.6

### v3

@@ 1 Introduction
Figure 1: Replay allows for matching repeated from-scratch training. Each line traces a continually trained 3B model where the annotations indicate the data cutoff year of the checkpoint (e.g., “24” = 2024). We find that combining autoregressive (AR) learning rate schedules and data replay (red) can nearly match the perplexity on all months achieved by the Oracle series which re-trains from scratch every two years (gray), despite requiring 2.6$\times$ less compute. Meanwhile, methods that only modify the optimizer (blue) or loss (green) insufficiently prevent forgetting and plateau.
@@ 3 TiC-CC: over 10 years of web data
Data processing. We build upon the assets from DataComp-LM (Li et al., 2024a), starting with DCLM-Pool (Li et al., 2024a), which contains all CC dumps up to Dec-2022 and parsed to extract plaintext from webpages via the resiliparse library (Bevendorff et al., 2018, 2021). We split this data by month and reuse the same download and processing scripts to extend DCLM-Pool until July-2024. Next, we follow DCLM-Baseline’s pipeline by applying heuristic filters from RefinedWeb (Penedo et al., 2023) and a fuzzy-deduplication step which we modify to run only within each month rather than across all months. Alternatively, as in TiC-CLIP, one could deduplicate newer months against older ones. Here, we do not do this by default given that it may not always be helpful (e.g., it may remove updated or edited pages where a few key facts have changed but most text remains the same). We perform an initial exploration of cross-month deduplication in Section D.7 and overall allow for exploring the benefits/pitfalls of such data-centric interventions as part of method design. Finally, we do not use the final classifier-based filter in DCLM-Baseline, as this classifier was trained on data from all months, violating causality. For more details about the data pipeline, see Appx. A.
@@ 5 Continual Learning Baseline Methods
  • Cyclic Cosine is the simplest alternative which applies cosine decay within each training month using the same maximum learning rate and warmup (Fig. 15, green). This was found to be most effective in TiC-CLIP (Garg et al., 2024).
  • Cyclic Cosine + AR (autoregressive) uses Cyclic Cosine but also decays the maximum learning rate across rounds, regressed from a global cosine schedule (Fig. 15, blue). It was shown to offer improvements by Roth et al. (2024).
@@
Non-continual Oracles. The alternative to continual learning is to simply retrain a model from scratch on an equal number of tokens from all currently available months. We refer to such a model as Oracle-$t$ where $t$ is particular cutoff date. We then consider as a baseline a series of Oracle models re-trained roughly every two years (i.e., {2013-05, 2015-01, 2017-01, 2019-01, 2021-01, 2023-01, 2024-07}), where for any timestamped evaluation, the most recently trained Oracle is always used.
@@ 6 Experiments
Figure 3: Continual methods incur different trade-offs between ID and Backward performance on TiC-CC. We plot selected regret matrices for 3B models and 440B training tokens (lower is better). Overall, Cyclic Cosine leads to strong ID performance (along the diagonal) but also significant forgetting after a few years. This can be partially addressed by using an AR meta-schedule and more significantly by replay. However, too much replay such as with $\alpha_{t}=1/t$ scales poorly with a large number of timestamps, significantly sacrificing ID performance.
Table 2: Loss-based evaluations for various methods. We report log-perplexity relative to Oracle-2024-07 models of the same size. While optimizer (top) and regularization-based (bottom) methods trade-off backward transfer with ID performance, replay (middle) is required to obtain the least amount of forgetting. Bold values are within one standard deviation of the best in each column for a given token budget, with standard deviations estimated from three runs of Cyclic Cosine. Highlighted values indicate when a continual run outperforms the Oracle series.
Method Tokens Params TiC-CC $\downarrow$ TiC-CC-Wiki $\downarrow$ TiC-CC-News $\downarrow$
@@
Cyclic Cosine (std) 220B 1B 0.079 0.026 0.154 0.041 0.033 0.072 0.065 0.016 0.105
Cyclic Cosine + AR 220B 1B 0.065 0.040 0.158 0.036 0.033 0.073 0.047 0.017 0.106
Cyclic Rsqrt 220B 1B 0.073 0.031 0.154 0.039 0.032 0.071 0.057 0.015 0.104
Schedule-Free 220B 1B 0.075 0.036 0.158 0.043 0.037 0.075 0.058 0.019 0.107
Replay ($\alpha=1/t$) 220B 1B 0.026 0.072 0.170 0.024 0.036 0.076 0.008 0.034 0.112
Replay ($\alpha=1/2$) 220B 1B 0.028 0.042 0.159 0.026 0.032 0.073 0.017 0.019 0.107
Replay ($\alpha=1/t$) + AR 220B 1B 0.029 0.080 0.173 0.022 0.038 0.076 0.007 0.038 0.113
Replay ($\alpha=1/2$) + AR 220B 1B 0.029 0.054 0.163 0.025 0.033 0.073 0.012 0.022 0.107
LwF 220B 1B 0.079 0.027 0.154 0.042 0.034 0.073 0.066 0.016 0.105
EWC 220B 1B 0.067 0.035 0.155 0.035 0.031 0.070 0.050 0.015 0.103
Oracle Series 1.16T 1B -0.003 0.035 0.154 0.004 0.016 0.062 -0.008 0.019 0.100
Cyclic Cosine (std) 220B 3B 0.072 0.027 0.161 0.038 0.032 0.074 0.058 0.015 0.109
Cyclic Cosine + AR 220B 3B 0.058 0.040 0.166 0.032 0.031 0.074 0.041 0.017 0.110
Cyclic Rsqrt 220B 3B 0.065 0.030 0.162 0.033 0.030 0.073 0.049 0.015 0.108
ScheduleFree 220B 3B 0.065 0.036 0.164 0.036 0.033 0.076 0.049 0.017 0.110
Replay ($\alpha=1/t$) 220B 3B 0.023 0.074 0.178 0.020 0.036 0.078 0.005 0.035 0.117
Replay ($\alpha_{t}=1/2$) 220B 3B 0.024 0.042 0.167 0.024 0.031 0.074 0.013 0.019 0.111
Replay ($\alpha=1/t$) + AR 220B 3B 0.026 0.083 0.181 0.019 0.037 0.079 0.004 0.039 0.119
Replay ($\alpha=1/2$) + AR 220B 3B 0.025 0.055 0.171 0.022 0.032 0.076 0.009 0.022 0.112
LwF 220B 3B 0.072 0.027 0.161 0.038 0.032 0.074 0.058 0.015 0.109
EWC 220B 3B 0.061 0.032 0.162 0.031 0.029 0.071 0.046 0.014 0.108
Cyclic Cosine (std) 440B 3B 0.082 -0.011 0.145 0.029 0.015 0.059 0.071 -0.001 0.099
Cyclic Cosine + AR 440B 3B 0.058 -0.002 0.148 0.014 0.009 0.057 0.044 -0.005 0.097
Cyclic Rsqrt 440B 3B 0.067 -0.007 0.146 0.018 0.010 0.057 0.055 -0.004 0.097
Schedule-Free 440B 3B 0.063 -0.004 0.147 0.017 0.011 0.059 0.049 -0.004 0.098
Replay ($\alpha=1/t$) 440B 3B 0.001 0.044 0.164 0.003 0.016 0.062 -0.009 0.013 0.105
Replay ($\alpha=1/2$) 440B 3B 0.007 0.007 0.151 0.010 0.013 0.059 0.005 -0.000 0.099
Replay ($\alpha=1/t$) + AR 440B 3B -0.003 0.050 0.166 -0.006 0.013 0.061 -0.017 0.014 0.105
Replay ($\alpha=1/2$) + AR 440B 3B -0.002 0.016 0.154 -0.001 0.009 0.057 -0.009 -0.002 0.098
LwF 440B 3B 0.082 -0.011 0.145 0.029 0.015 0.059 0.072 -0.001 0.099
EWC 440B 3B 0.055 0.011 0.152 0.017 0.015 0.061 0.041 0.001 0.100
Oracle Series 1.16T 3B -0.003 0.037 0.163 0.004 0.017 0.066 -0.007 0.020 0.107
@@
Training details. We train 1B and 3B parameter models using OpenLM (Gururangan et al., 2023). Unless otherwise indicated, most of our results are for 3B models, where each method sees the same total number of 220B or 440B tokens, equivalent to 4$\times$ and 8$\times$ the Chinchilla optimal amount.  [^2] Here, token counts are given by 20 $\times$ parameters $\times$ Chinchilla multiplier with a 1$\times$ multiplier being a near-optimal compute allocation found by Hoffmann et al. (2022). For 1B models, we only train at the 220B token scale to study how model size might affect the results (see Table 2 and Section D.8). For continual runs, we further assume that current practitioners are (a) likely to have access to more than enough data to train initial models; (b) unlikely to wait to obtain non-trivial performance. Hence, we front-load the total token budget such that 110B is allocated to an initial pretraining on the first month (May-2013). Then, the remaining tokens are split equally among the other 113 continual timesteps. For our Oracle-$t$ runs, each is trained on the same number of tokens that a 220B continual run observes by the end of month $t$, totaling 1.16T tokens for all seven Oracle models together. Finally, to perform realistic hyperparameter selection, we follow takeaways from (Cha and Cho, 2024) and only use the first 10 timesteps for tuning (see Appx. C for more details).
@@
Evaluation metrics. Each run produces a $T_{t}\times T_{e}$ matrix of evaluations $E$ where $T_{t},T_{e}$ are the total number of training/evaluation timesteps, $E_{i,j}$ is the performance of the model trained after seeing data up to month $i$ and evaluated on the month $j$. To control for inherent difficulty differences across evaluation months, we measure the regret $R_{i,j}=E_{i,j}-E^{*}_{j}$ where $E^{*}_{j}$ is the performance of Oracle-2024-07 on month $j$. Following Garg et al. (2024), we consider the following summary metrics, first defined assuming $T_{t}=T_{e}=T$ (deferring the discussion of the $T_{t}\neq T_{e}$ case to Sec. B.4):
@@
  • Backward transfer: averages the lower triangular of $R$, i.e., $\sum_{i=1}^{T}\sum_{j<i}\frac{R_{i,j}}{T(T-1)/2}$, capturing how well continual checkpoints do on older months.
  • Forward transfer: averages the upper triangular of $R$ analogously to backward transfer, capturing how well methods do on unseen future months.
@@
6.1 Held-out performance on TiC-CC
Table 3: Benchmarking continual methods on downstream evaluations (3B Models). For all dynamic evaluations, we report perplexity relative to Oracle-2024-07 with log-scaling. Core is an average of the accuracies of 22 downstream zero/few-shot tasks used by Li et al. (2024a), evaluated only on the final model checkpoint (without scaling by an Oracle). Bold values are within one standard deviation (estimated with 3 runs of Cyclic Cosine) of the best in each column. Highlighted values indicate matching the Oracle series (within one standard deviation).
[…]
Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd. Bwd. ID Fwd.
Cyclic Cosine (std) 440B 0.018 0.026 0.072 0.026 0.031 0.056 0.017 0.045 0.142 -0.020 -0.002 0.175
Cyclic Cosine + AR 440B 0.010 0.023 0.071 0.010 0.022 0.054 0.026 0.057 0.142 0.012 0.027 0.195
Cyclic Rsqrt 440B 0.012 0.024 0.070 0.014 0.024 0.053 0.024 0.055 0.145 -0.001 0.015 0.187
ScheduleFree 440B 0.012 0.025 0.073 0.014 0.025 0.056 0.038 0.066 0.144 0.030 0.044 0.197
Replay ($\alpha=1/t$) 440B 0.019 0.038 0.078 0.014 0.028 0.057 0.039 0.087 0.176 0.106 0.119 0.238
Replay ($\alpha=1/2$) 440B 0.015 0.029 0.073 0.017 0.027 0.055 0.027 0.065 0.158 0.020 0.032 0.189
Replay ($\alpha=1/t$) + AR 440B 0.015 0.036 0.077 0.005 0.023 0.056 0.057 0.102 0.179 0.141 0.151 0.261
Replay ($\alpha=1/2$) + AR 440B 0.010 0.027 0.073 0.007 0.022 0.054 0.036 0.074 0.160 0.058 0.070 0.214
LwF 440B 0.019 0.027 0.072 0.026 0.031 0.056 0.014 0.047 0.142 -0.028 -0.011 0.171
EWC 440B 0.015 0.030 0.074 0.016 0.029 0.058 0.039 0.073 0.155 0.055 0.069 0.211
Oracle Series 1.16T 0.014 0.035 0.080 0.013 0.030 0.061 0.012 0.056 0.146 0.035 0.057 0.196
Bwd. ID Fwd. Bwd. ID Fwd. Core (DCLM)
Cyclic Cosine 440B 0.048 0.030 0.013 0.054 0.070 0.065 49.6
Cyclic Cosine + AR 440B 0.023 0.008 -0.001 0.044 0.057 0.055 48.8
Cyclic Rsqrt 440B 0.032 0.017 0.006 0.050 0.061 0.060 49.4
ScheduleFree 440B 0.036 0.023 0.010 0.081 0.091 0.072 49.4
Replay ($\alpha=1/t$) 440B -0.025 -0.025 -0.022 0.035 0.038 0.041 49.4
Replay ($\alpha=1/2$) 440B $0$0.008 -0.011 0.029 0.038 0.044 50.1
Replay ($\alpha=1/t$) + AR 440B -0.028 -0.027 -0.022 0.031 0.032 0.047 49.3
Replay ($\alpha=1/2$) + AR 440B -0.015 -0.019 -0.017 0.036 0.039 0.043 49.3
LwF 440B 0.049 0.031 0.013 0.059 0.077 0.070 49.6
EWC 440B 0.014 0.004 0.000 0.063 0.066 0.059 49.0
Oracle series 1.16T -0.025 -0.028 -0.022 0.008 0.008 0.015 50.6
In Tab. 2 and Fig. 3, we first explore how well continual methods learn and retain various parts of our general web-scale training distribution, TiC-CC. Overall, we observe significant trade-offs between ID and backward transfer, with key findings below:
@@ 6.1 Held-out performance on TiC-CC
Replay is essential for addressing forgetting. Based on backward transfer and Figs. 17 and 16, all non-replay methods show significant forgetting at later checkpoints. Optimizer tweaks and EWC can somewhat reduce forgetting by sacrificing some ID performance. However, additionally applying replay can further improve backward transfer by 60% for 220B runs, a gap which only widens as we scale up to 440B tokens. Between replay variants, $\alpha_{t}=1/t$ achieves the lowest forgetting but $\alpha_{t}=1/2$ offers a better practical trade-off on TiC-CC, resulting in marginally worse Backward but substantially better ID. This differs from TiC-CLIP’s recommendation of $\alpha_{t}=1/t$, likely since decreasing the ratio of new data becomes problematic in our setup which uses 10$\times$ more timesteps.
@@
For smaller models, forgetting is more of a challenge but method rankings remain similar. We observe that ID for 1B methods (relative to the 1B Oracle) is quite close to ID for 3B methods (relative to the 3B Oracle). However, Backward metrics are worse at 1B especially when not using replay. This suggests that, perhaps as expected, models struggle more to retain older knowledge when they have fewer parameters. Overall though, we obtain similar conclusions between the two scales in terms of what methods best mitigate forgetting: increasing model size can slightly improve relative Backward regret but not nearly as much as switching from non-replay to replay methods.
@@ 6.2 Downstream evaluations
Continual methods outperform Oracle re-training on many downstream evaluations. While Replay ($\alpha_{t}=1/2$) + AR at 440B tokens outperforms the Oracle series on TiC-CC, the comparison is more nuanced for downstream tasks. Many continual methods show significant gains on TiC-Wiki evaluations, while the Oracle series maintains an edge on TiC-CodeDocs-NumPy and TiC-StackE-English (see Appx. D). Meanwhile, methods like Cyclic Cosine outperform the Oracles on TiC-StackOverflow and TiC-CodeDocs-PyTorch but fall short on TiC-StackE-Math where only Replay $(\alpha=1/t)$ can do so.